OpenStudy (anonymous):

Radical Equation: (x^2-4)1/2 =4

OpenStudy (anonymous):

idk if thats right

OpenStudy (anonymous):

I need step by step process.. giving me the answer is not helping.

OpenStudy (anonymous):

multiply by 2 on both sides

OpenStudy (anonymous):

add 4 to both sides

OpenStudy (anonymous):

square both sides

OpenStudy (anonymous):

Aren't you suppose to multiply 1/2 to both sides?

OpenStudy (anonymous):

could you rewrite your prob using the equation button below

OpenStudy (anonymous):

\[(x ^{2}-4)1/2 =4\]

OpenStudy (anonymous):

is 1/2 separate from (x^2-4)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so then divide by 1/2 om both sides

OpenStudy (anonymous):

so u get x^2-4=2

OpenStudy (anonymous):

then add 4

OpenStudy (anonymous):

so u get x^2=6

OpenStudy (anonymous):

then square both sides

OpenStudy (anonymous):

x=\[\sqrt{6}\]

OpenStudy (anonymous):

did i do it wrong?